OpenStudy (anonymous):

WILL MEDAL!!!! 6. Suppose that the diameters of different types of hairs are 1.7*10^-4cm, 1.4*10^-3cm, 1.2*10^-5cm, and 1.9*10^-4cm? What is the order of the hairs from least to greatest diameter? a. 1.9*10^-4, 1.7*10^-4, 1.4*10^-3, 1.2*10^-5 b. 1.2*10^-5, 1.7*10^-4, 1.9*10^-4, 1.4*10^-3 c. 1.9*10^-4, 1.7*10^-4, 1.2*10^-5, 1.4*10^-3 d. 1.2*10^-5, 1.4*10^-3, 1.9*10^-4, 1.7*10^-4 7. What is the expression written using each base only once? (-4)^-6 * (-4)^7 a. 4^13 b. -4 c. 13 d. 1

OpenStudy (anonymous):

@thedogkeeper

OpenStudy (anonymous):

1,47*10^-4=1.47/1000 so as the denominator get bigger the value get smaller .... 10^-5 is the smallest

OpenStudy (anonymous):

-4^-6*-4^7 during multiplication we add the powers if the base are the same -4^(-6+7)=-4^1=-4

OpenStudy (anonymous):

the answer for both is B